# Modern Neuromorphic AI: From Intra-Token to Inter-Token Processing

### v1

@@ I Introduction
Modern artificial intelligence (AI) tools are widely believed to have the potential to transform our economies and societies, yet this progress comes at a well-documented steep cost in terms of energy resources for both training [cottier2024rising] and inference [smith2025aienergy]. The quest for a large “intelligence-to-joule ratio” [sambanova2025intelligence] has long been the objective of neuromorphic engineering, which takes inspiration from the most energy-efficient intelligent system known: the human brain [mead2020we]. Over more than three decades of research, neuromorphic engineering has identified a number of principles from neuroscience that may be responsible for the energy efficiency of the brain [davies2021advancing, rajendran2025], including: asynchronous event-driven information acquisition, sparse communication through discrete spikes (i.e., binary signals), dynamic state-space computations in neuronal membranes, analog in-memory computing, massively parallel computing, and local learning rules.
@@
Recent developments in mainstream AI have incorporated a number of lessons from neuromorphic engineering [rajendran2025]. Most notably, state-of-the-art large language models (LLMs) have been demonstrated with activations quantized to integers with resolutions as low as 4 bits [ma2024era, liu2023llm] – moving towards the neuromorphic regime of spike-based (i.e., binary) communication. Furthermore, some of the best-performing sequence-to-sequence models are designed based on state dynamics – reproducing the stateful operation of typical neuromorphic models [gu2021efficiently, gu2023mamba]. This convergence reflects fundamental computational principles: efficient systems must exploit temporal structure if present and avoid unnecessary computation through sparsity.
Despite the outlined convergence of AI along neuromorphic principles, the research on neuromorphic AI has remained relatively separate from mainstream AI. The state of the art up to around 2022 is well summarized in a number of tutorial papers, including the November 2019 special issue of IEEE Signal Processing Magazine on “Learning Algorithms and Signal Processing for Brain-Inspired Computing” [neftci2019surrogate, jang2019introduction, indiveri2019importance, rajendran2019low], and later reviews on event-driven sensing [chen2020event] and neuroscience-inspired learning algorithms [pehlevan2019neuroscience]. Other surveys have organized the field around algorithm-hardware co-design for spike-based machine intelligence [roy2019towards], and specific platforms such as Intel’s Loihi processor [davies2021advancing]. These reviews have established neuromorphic computing as a mature field with demonstrated applications in vision, robotics, and signal processing.
@@
Inter-token processing (Figure 5) mixes information between different tokens in a sequence, separately for each channel. Notably, in transformers, self-attention mechanisms implement inter-token processing by computing pairwise relationships between all tokens; while in SSMs, state-space dynamic updates carry information from past to future tokens. Recent work has explored the adoption of neuromorphic principles for the implementation of inter-token processing, encompassing both self-attention [zhou2023spikformer, song2025xpikeformer, xing2024spikellm, pan2025spikingbrain] and SSMs [stan2024learning, bal2024pspikessm]. In such systems, as illustrated in the bottom part of Figure 4 and exemplified in Table II, improved energy-accuracy trade-offs arise from the sparsity of the inner representations leveraged by neuromorphic AI systems.
@@ II-A Tokenization and Embedding
Consider the case of an input text. As seen in Figure 1, the text is first partitioned into discrete word or subword units using algorithms like byte pair encoding [sennrich2016neural], converting it into a sequence of token identifiers (IDs). This process is known as tokenization. Each token ID is an integer from the set of integers $\{1,...,V\}$, where $V$ is the size of the dictionary of possible tokens.
@@
Each token ID $v\in\{1,...,V\}$ is then mapped to a real-valued vector of dimension $D_{\text{emb}}$}, which is read off a fixed dictionary of $V$ vectors. The vector assigned to each token ID is known as token embedding. Each entry of the embedding vector is known as a channel (or feature) of the embedding. Embeddings may also be pre-processed to include positional information about the index $n$, since the original embeddings alone do not carry this information. This is typically done by summing vectors or by multiplying by matrices that depend on index $n$ [su2024roformer, dufter2022position].
@@
Similar transformations can also be applied to other types of inputs, including images and videos. To this end, an image is first divided into fixed-size patches, which are flattened and multiplied by a matrix to produce patch embeddings [dosovitskiy2021image]. In this case, the token index $n$ does not represent a temporal index, but is rather an indicator for the spatial position of the token in the image. Tokens are typically pre-processed to encoded this information. For videos, frames are sampled temporally, and each frame is processed like an image, with additional temporal positional information added to capture the sequential nature of frames [arnab2021vivit, bertasius2021spacetime].
@@ II-B Modern AI Architectures
Keyword spotting $109\times$ [blouw2019benchmarking, Figure 1]
CIFAR-10 classification $100\times$ [lenz2023ultra, Figure 4]
Image segmentation $2\times$ [patel2021spiking, Figure 12]
@@
Language modeling RWKV-based SNN $32.2\times$ [zhu2023spikegpt, Table 1]
Language modeling Self-attention (BERT-based) $2\times$ [bal2024spikingbert, Figure 4]
Long Range Arena SSM-based SNN $36\times$ [bal2024pspikessm, Sec. 4.2]
Symbol demodulation Self-attention (no masking) $14\times$ [song2025xpikeformer, Figure 8]
@@ II-E From Intra-Token to Inter-Token Processing
In such applications, the temporal dynamics of neuromorphic models is leveraged to carry out inter-channel mixing over a virtual time axis (see Figure 7). This may yield energy savings as compared to standard neural networks when the signals processed over this virtual time dimension are sparse and outputs can be produced in a few steps of the virtual time axis [davies2021advancing] (see Figure 4). In fact, energy consumption grows with the number of virtual time steps, while decreasing with the sparsity of the input. Examples of representative results concerning the energy advantages of SNNs implemented over specialized hardware against conventional neural networks run of GPUs can be found in Table I.
@@
Conventional examples of neuromorphic models processing information over real time encompass SNNs deployed to extract information from event-driven data produced by neuromorphic sensors [schaefer2022aegnn, anumula2018feature, chen2023neuromorphic]. As discussed later in this article, more recent examples include spiking language models, in which discrete time runs over embeddings of a given text [zhu2023spikegpt, yao2023spikedriven, xing2024spikellm, pan2025spikingbrain] and token-based vision models [song2025xpikeformer]. Table II provides some examples of existing experimental findings about the potential advantages of neuromorphic computing for tasks requiring inter-token processing.
@@ III-A Defining Neuromorphic Processing Elements
    • Dynamic sparsity: Downstream NPEs can avoid processing zero activations, since zero-valued inputs do not change their outputs. Therefore, a larger sparsity in the signals exchanged by NPEs – in virtual or real time – can yields important gains in terms of energy consumption, as long as the hardware makes it possible to adaptively gate unnecessary operations [rathi2023exploring].
    • Multiplication-free operations: When the activations are binary, multiplications with weight matrices require only accumulate (sum) operations, hence foregoing the need for significantly more complex multiply-and-accumulate operations. In fact, a multiplier typically consumes several times the energy of an adder, while also requiring an orders-of-magnitude larger chip area [horowitz20141].
@@ III-B2 Internal State
The LIF model maintains a real-valued state vector $\bm{u}[t]$, which is updated via a scalar gain $\alpha$. Alternatively, there exist spiking neural models that implement higher-order dynamics by introducing more than one state variable per output $y_{i}[t]$ and by applying matrix gains in the state update. This approach enables richer temporal dynamics, including oscillatory behavior. Accordingly, these models are particularly well suited to model periodic patterns and, more broadly, signals with informative spectral features, such as audio or radio signals [izhikevich2001resonate, higuchi2024balanced, wu2025neuromorphic]. As an alternative to state-based dynamics of the form (9), transition models based on look-up tables operating at the level of groups of timings have also been proposed [izhikevich2025spikingmanifesto].
@@ IV-A Input Encoding Along the Virtual Time Axis
Given a real-valued input token $\bm{X}_{:,n}$, the first step in the deployment of an NPU is the encoding of the vector $\bm{X}_{:,n}$ as a signal that evolves along the virtual time axis $t$. This conversion is typically necessary at the input of the overall NPU, i.e., at the bottom of the architecture in Figure 2. However, in some applications, the input to the NPU is already in the form of spiking signals. This is the case for NPUs that process inputs from neuromorphic sensors, such as as event-driven cameras, that produce spiking signals [chen2020event].
@@ IV-A3 Time Encoding
Rate encoding is information-theoretically inefficient, as it encodes $\log_{2}(T+1)$ bits in $T$ virtual time steps. In fact, there are only $T+1$ values that the rate (11) can take. Other, potentially more efficient, encoding strategies include time encoding, in which the different virtual times $t\in\{1,...,T\}$ play a distinct role. In particular, with positional time encoding, the $T$ bits $x_{d}[1],...,x_{d}[T]$ represent a $T$-bit integer, thus carrying up to $T$ bits of information. More constrained forms of time encoding, e.g., based on first-to-spike time encoding or relative-timing encoding, have a generally lower capacity [jang2019introduction, izhikevich2025spikingmanifesto] (see the top-right panel of Figure 7 for an illustration of first-to-spike encoding).
@@ IV-A4 Multi-level spikes
The virtual time horizon $T$ can be potentially compressed by using multi-level spikes instead of binary spikes. For example, assuming rate encoding, each $L$-ary spike can encode the information contained in $L$ virtual time steps [xing2024spikellm] (see Figure 7 for $L=2$). As another example, with positional time encoding, each $L$-ary spike can encode $\log_{2}(L+1)$ time steps.
@@ V-A Inter-Token Mixing as Associative Memorization
The intra-token mixing steps applied by both SSMs and transformers can be viewed as forms of associative memorization. Associative memories implement content-addressable information retrieval mechanisms. Unlike conventional memories, the output is obtained not based on a specific address, but by matching cue to stored patterns based on content similarity [krotov2025modern].
@@
As a side note, it is also possible to formalize intra-token processing in two-layer feedforward neural networks, which are typically adopted in transformer architectures, as a form of associative memory. However, unlike the dynamic and adaptive associative memory mechanism applied by inter-token mixing, intra-token processing effectively implements a static memorization mechanism [geva2021transformer].
@@ V-B Inter-Token Processing via State-Space Models
To this end, for each $d$-th channel, SSMs maintain a vector state $\bm{h}_{d}[t]\in\mathbb{R}^{D_{\text{k}}}$ of the same dimensions, $D_{\text{k}}$, of the key and query vectors as [lahoti2025chimera]
@@
  • Gating: The gains $a[t]$ and $b[t]$ in the state-space dynamics (16) can be adapted over time $t$ to implement gating mechanisms. Gating modulates the extent to which input $\bm{k}[t]v_{d}[t]$ is memorized via the state $\bm{h}[t]$. This yields the class of selective SSMs. A common choice is $a[t]=\exp(-\Delta[t])$ and $b[t]=\Delta[t]$, where $\Delta[t]$ is a non-negative function of the input $\bm{Z}_{:,t}$ [lahoti2025chimera]. This way, a larger value of the time step $\Delta[t]$ causes the value channel $v_{d}[t]$ to affect the state $\bm{h}_{d}[t]$ to a larger extent, partially “forgetting” past values. A similar mechanism can be potentially implemented also for intra-channel mixing [behrouz2024mambamixer].
  • Matrix-based gains: Instead of using a scalar gain $a$, it is possible to multiply the past value vector $\bm{h}[t-1]$ with a matrix $\bm{A}$. This matrix can be designed to approximate dynamic projections into an evolving orthogonal basis [gu2022hippo].
@@ V-C NPEs for Inter-Token Processing via State-Space Dynamics
Beyond the presence of the reset mechanism, the LIF-based NPE, however, deviates from an SSM (16)-(17) due to the adoption of non-linear activation functions (4)-(7). Stacking NPEs with reset-free dynamics and with the probabilistic spiking mechanism (5) yields the P-SpikeSSM architecture introduced in [bal2024pspikessm] (see also [stan2024learning]).
@@ V-D RWKV
Inter-token processing based on state-space dynamics of the form (16) leverages a state vector to memorize past key-value associations. This memory is then used to retrieve the new token value by projecting the state onto the new query vector. In contrast, the receptance weighted key value model (RWKV) implements a form of inter-token processing by directly mixing value vectors element-wise before producing the output via a gating mechanism followed by a linear projection [peng2023rwkv]. Value mixing takes place in a recurrent way, implementing a form of state-space dynamics [peng2023rwkv, Appendix D].
@@
In a manner similar to the relationship between SSMs and P-SpikeSSM, SpikeGPT adapts RWKV for neuromorphic computing by following the described RWKV processing with a layer of LIF neurons producing binary activations [zhu2023spikegpt].
@@ VI-B Attention-Based NPEs
As discussed, softmax self-attention require the evaluation of the query-key inner products $\bm{Q}_{n,:}^{\top}\bm{K}_{n^{\prime},:}$ in the attention scores (23), as well as the attention-value inner products in (22). These inner products generally require the implementation of a number of multiply-and-accumulate operations that scales with the size of the vectors. Alternatively, as proposed in recent research, approximations of these inner products can be evaluated more efficiently by using encoding of the queries, keys, and values as sparse spiking signals along a virtual time axis. This approach forms the core of a number of spiking transformers [zhou2023spikformer, zhou2023spikingformer, abreu2025neuromorphic, song2025xpikeformer]. As illustrated in Figure 8, the following main strategies have been proposed, which operate on binary spiking signals.
  • LIF-based encoding with coincidence-based scores: The linear projections (14) are first converted into spiking signals by passing each vector through a LIF-based NPE. As a result, one obtains a spiking signal, defined along the virtual time axis $t$, per vector. The similarity between the two vectors is then evaluated via a direct comparison of the corresponding binary vectors. To this end, a natural idea is to count the number of coincident spikes in the two vectors. This requires only a logical AND operation per entry to detect two spikes in the same position [zhou2023spikformer, zhou2023spikingformer, abreu2025neuromorphic]. This metric is extremely cheap to evaluate, and it intuitively captures a notion of similarity encoded in the timings of the spikes in the two vectors. An alternative metric counts the number of coincident binary digits – thus including not only ones but also zeros. This metric requires an XNOR operation per entry of the encoded time series [xiao2025rethinking]. Figure 8 illustrates an example for spiking encodings corresponding to a key vector and a query vector.
  • Stochastic computing: The work [song2025xpikeformer] leverages stochastic computing [alaghi2013survey] to evaluate the inner products required by self-attention. Unlike the LIF-based encoding schemes used by the previous class of methods, stochastic computing recovers exact results as the virtual time index $t$ grows large. To this end, queries, keys, and values are transformed into binary signals via Bernoulli encoding (recall (12)). This representation enables multiplications to be performed using simple AND gates. In fact, the result of the AND of two independent Bernoulli random variables $\text{Bern}(p_{1})$ and $\text{Bern}(p_{2})$ is a Bernoulli random variable $\text{Bern}(p_{1}p_{2})$. Therefore, by the law of large numbers, as the virtual time index increases, the estimated rate of spike coincidences detected by the AND gates tends to the correct product value $p_{1}p_{2}$.

### v2

@@ I Introduction
Modern artificial intelligence (AI) tools are widely believed to have the potential to transform our economies and societies, yet this progress comes at a well-documented steep cost in terms of energy resources for both training [15] and inference [54]. The quest for a large “intelligence-to-joule ratio” [50] has long been the objective of neuromorphic engineering, which takes inspiration from the most energy-efficient intelligent system known: the human brain [40]. Over more than three decades of research, neuromorphic engineering has identified a number of principles from neuroscience that may be responsible for the energy efficiency of the brain [17, 47], including: asynchronous event-driven information acquisition, sparse communication through discrete spikes (i.e., binary signals), dynamic state-space computations in neuronal membranes, analog in-memory computing, massively parallel computing, and local learning rules.
@@
Recent developments in mainstream AI have incorporated a number of lessons from neuromorphic engineering [47]. Most notably, state-of-the-art large language models (LLMs) have been demonstrated with activations quantized to integers with resolutions as low as 4 bits [39, 38] – moving towards the neuromorphic regime of spike-based (i.e., binary) communication. Furthermore, some of the best-performing sequence-to-sequence models are designed based on state dynamics – reproducing the stateful operation of typical neuromorphic models [26, 25]. This convergence reflects fundamental computational principles: efficient systems must exploit temporal structure if present and avoid unnecessary computation through sparsity.
Despite the outlined convergence of AI along neuromorphic principles, the research on neuromorphic AI has remained relatively separate from mainstream AI. The state of the art up to around 2022 is well summarized in a number of tutorial papers, including the November 2019 special issue of IEEE Signal Processing Magazine on “Learning Algorithms and Signal Processing for Brain-Inspired Computing” [41, 34, 31, 46], and later reviews on event-driven sensing [13] and neuroscience-inspired learning algorithms [44]. Other surveys have organized the field around algorithm-hardware co-design for spike-based machine intelligence [49], and specific platforms such as Intel’s Loihi processor [17]. These reviews have established neuromorphic computing as a mature field with demonstrated applications in vision, robotics, and signal processing.
@@
Inter-token processing (Figure 5) mixes information between different tokens in a sequence, separately for each channel. Notably, in transformers, self-attention mechanisms implement inter-token processing by computing pairwise relationships between all tokens; while in SSMs, state-space dynamic updates carry information from past to future tokens. Recent work has explored the adoption of neuromorphic principles for the implementation of inter-token processing, encompassing both self-attention [67, 55, 62, 42] and SSMs [56, 5, 65]. In such systems, as illustrated in the bottom part of Figure 4 and exemplified in Table II, improved energy-accuracy trade-offs arise from the sparsity of the inner representations leveraged by neuromorphic AI systems.
@@ II-A Tokenization and Embedding
Consider the case of an input text. As seen in Figure 1, the text is first partitioned into discrete word or subword units using algorithms like byte pair encoding [52], converting it into a sequence of token identifiers (IDs). This process is known as tokenization. Each token ID is an integer from the set of integers $\{1,...,V\}$, where $V$ is the size of the dictionary of possible tokens.
@@
Each token ID $v\in\{1,...,V\}$ is then mapped to a real-valued vector of dimension $D_{\text{emb}}$}, which is read off a fixed dictionary of $V$ vectors. The vector assigned to each token ID is known as token embedding. Each entry of the embedding vector is known as a channel (or feature) of the embedding. Embeddings may also be pre-processed to include positional information about the index $n$, since the original embeddings alone do not carry this information. This is typically done by summing vectors or by multiplying by matrices that depend on index $n$ [57, 20].
@@
Similar transformations can also be applied to other types of inputs, including images and videos. To this end, an image is first divided into fixed-size patches, which are flattened and multiplied by a matrix to produce patch embeddings [18]. In this case, the token index $n$ does not represent a temporal index, but is rather an indicator for the spatial position of the token in the image. Tokens are typically pre-processed to encoded this information. For videos, frames are sampled temporally, and each frame is processed like an image, with additional temporal positional information added to capture the sequential nature of frames [4, 10].
@@ II-B Modern AI Architectures
Keyword spotting $109\times$ [12, Figure 1]
CIFAR-10 classification $100\times$ [37, Figure 4]
Image segmentation $2\times$ [43, Figure 12]
@@
Language modeling RWKV-based SNN $32.2\times$ [68, Table 1]
Language modeling Self-attention (BERT-based) $2\times$ [6, Figure 4]
Long Range Arena SSM-based SNN $36\times$ [5, Sec. 4.2]
Symbol demodulation Self-attention (no masking) $14\times$ [55, Figure 8]
@@ II-E From Intra-Token to Inter-Token Processing
In such applications, the temporal dynamics of neuromorphic models is leveraged to carry out inter-channel mixing over a virtual time axis (see Figure 7). This may yield energy savings as compared to standard neural networks when the signals processed over this virtual time dimension are sparse and outputs can be produced in a few steps of the virtual time axis [17] (see Figure 4). In fact, energy consumption grows with the number of virtual time steps, while decreasing with the sparsity of the input. Examples of representative results concerning the energy advantages of SNNs implemented over specialized hardware against conventional neural networks run of GPUs can be found in Table I.
@@
Conventional examples of neuromorphic models processing information over real time encompass SNNs deployed to extract information from event-driven data produced by neuromorphic sensors [51, 3, 14]. As discussed later in this article, more recent examples include spiking language models, in which discrete time runs over embeddings of a given text [68, 63, 62, 42] and token-based vision models [55]. Table II provides some examples of existing experimental findings about the potential advantages of neuromorphic computing for tasks requiring inter-token processing.
@@ III-A Defining Neuromorphic Processing Elements
    • Dynamic sparsity: Downstream NPEs can avoid processing zero activations, since zero-valued inputs do not change their outputs. Therefore, a larger sparsity in the signals exchanged by NPEs – in virtual or real time – can yields important gains in terms of energy consumption, as long as the hardware makes it possible to adaptively gate unnecessary operations [48].
    • Multiplication-free operations: When the activations are binary, multiplications with weight matrices require only accumulate (sum) operations, hence foregoing the need for significantly more complex multiply-and-accumulate operations. In fact, a multiplier typically consumes several times the energy of an adder, while also requiring an orders-of-magnitude larger chip area [29].
@@ III-B2 Internal State
The LIF model maintains a real-valued state vector $\bm{u}[t]$, which is updated via a scalar gain $\alpha$. Alternatively, there exist spiking neural models that implement higher-order dynamics by introducing more than one state variable per output $y_{i}[t]$ and by applying matrix gains in the state update. This approach enables richer temporal dynamics, including oscillatory behavior. Accordingly, these models are particularly well suited to model periodic patterns and, more broadly, signals with informative spectral features, such as audio or radio signals [32, 28, 60]. As an alternative to state-based dynamics of the form (9), transition models based on look-up tables operating at the level of groups of timings have also been proposed [33].
@@ IV-A Input Encoding Along the Virtual Time Axis
Given a real-valued input token $\bm{X}_{:,n}$, the first step in the deployment of an NPU is the encoding of the vector $\bm{X}_{:,n}$ as a signal that evolves along the virtual time axis $t$. This conversion is typically necessary at the input of the overall NPU, i.e., at the bottom of the architecture in Figure 2. However, in some applications, the input to the NPU is already in the form of spiking signals. This is the case for NPUs that process inputs from neuromorphic sensors, such as as event-driven cameras, that produce spiking signals [13].
@@ IV-A3 Time Encoding
Rate encoding is information-theoretically inefficient, as it encodes $\log_{2}(T+1)$ bits in $T$ virtual time steps. In fact, there are only $T+1$ values that the rate (11) can take. Other, potentially more efficient, encoding strategies include time encoding, in which the different virtual times $t\in\{1,...,T\}$ play a distinct role. In particular, with positional time encoding, the $T$ bits $x_{d}[1],...,x_{d}[T]$ represent a $T$-bit integer, thus carrying up to $T$ bits of information. More constrained forms of time encoding, e.g., based on first-to-spike time encoding or relative-timing encoding, have a generally lower capacity [34, 33] (see the top-right panel of Figure 7 for an illustration of first-to-spike encoding).
@@ IV-A4 Multi-level spikes
The virtual time horizon $T$ can be potentially compressed by using multi-level spikes instead of binary spikes. For example, assuming rate encoding, each $L$-ary spike can encode the information contained in $L$ virtual time steps [62] (see Figure 7 for $L=2$). As another example, with positional time encoding, each $L$-ary spike can encode $\log_{2}(L+1)$ time steps.
@@ V-A Inter-Token Mixing as Associative Memorization
The intra-token mixing steps applied by both SSMs and transformers can be viewed as forms of associative memorization. Associative memories implement content-addressable information retrieval mechanisms. Unlike conventional memories, the output is obtained not based on a specific address, but by matching cue to stored patterns based on content similarity [35].
@@
As a side note, it is also possible to formalize intra-token processing in two-layer feedforward neural networks, which are typically adopted in transformer architectures, as a form of associative memory. However, unlike the dynamic and adaptive associative memory mechanism applied by inter-token mixing, intra-token processing effectively implements a static memorization mechanism [24].
@@ V-B Inter-Token Processing via State-Space Models
To this end, for each $d$-th channel, SSMs maintain a vector state $\bm{h}_{d}[t]\in\mathbb{R}^{D_{\text{k}}}$ of the same dimensions, $D_{\text{k}}$, of the key and query vectors as [36]
@@
  • Gating: The gains $a[t]$ and $b[t]$ in the state-space dynamics (16) can be adapted over time $t$ to implement gating mechanisms. Gating modulates the extent to which input $\bm{k}[t]v_{d}[t]$ is memorized via the state $\bm{h}[t]$. This yields the class of selective SSMs. A common choice is $a[t]=\exp(-\Delta[t])$ and $b[t]=\Delta[t]$, where $\Delta[t]$ is a non-negative function of the input $\bm{Z}_{:,t}$ [36]. This way, a larger value of the time step $\Delta[t]$ causes the value channel $v_{d}[t]$ to affect the state $\bm{h}_{d}[t]$ to a larger extent, partially “forgetting” past values. A similar mechanism can be potentially implemented also for intra-channel mixing [7].
  • Matrix-based gains: Instead of using a scalar gain $a$, it is possible to multiply the past value vector $\bm{h}[t-1]$ with a matrix $\bm{A}$. This matrix can be designed to approximate dynamic projections into an evolving orthogonal basis [27].
@@ V-C NPEs for Inter-Token Processing via State-Space Dynamics
Beyond the presence of the reset mechanism, the LIF-based NPE, however, deviates from an SSM (16)-(17) due to the adoption of non-linear activation functions (4)-(7). Stacking NPEs with reset-free dynamics and with the probabilistic spiking mechanism (5) yields the P-SpikeSSM architecture introduced in [5] (see also [56] for a similar architecture). Neuromorphic models that retain the reset mechanism were studied in [19, 65]. There are also spiking SSMs that implement linear projections via intra-token NPEs along the virtual time axis [30].
@@ V-D RWKV
Inter-token processing based on state-space dynamics of the form (16) leverages a state vector to memorize past key-value associations. This memory is then used to retrieve the new token value by projecting the state onto the new query vector. In contrast, the receptance weighted key value model (RWKV) implements a form of inter-token processing by directly mixing value vectors element-wise before producing the output via a gating mechanism followed by a linear projection [45]. Value mixing takes place in a recurrent way, implementing a form of state-space dynamics [45, Appendix D].
@@
In a manner similar to the relationship between SSMs and P-SpikeSSM, SpikeGPT adapts RWKV for neuromorphic computing by following the described RWKV processing with a layer of LIF neurons producing binary activations [68].
@@ VI-B Attention-Based NPEs
As discussed, softmax self-attention require the evaluation of the query-key inner products $\bm{Q}_{n,:}^{\top}\bm{K}_{n^{\prime},:}$ in the attention scores (23), as well as the attention-value inner products in (22). These inner products generally require the implementation of a number of multiply-and-accumulate operations that scales with the size of the vectors. Alternatively, as proposed in recent research, approximations of these inner products can be evaluated more efficiently by using encoding of the queries, keys, and values as sparse spiking signals along a virtual time axis. This approach forms the core of a number of spiking transformers [67, 66, 1, 55]. As illustrated in Figure 8, the following main strategies have been proposed, which operate on binary spiking signals.
  • LIF-based encoding with coincidence-based scores: The linear projections (14) are first converted into spiking signals by passing each vector through a LIF-based NPE. As a result, one obtains a spiking signal, defined along the virtual time axis $t$, per vector. The similarity between the two vectors is then evaluated via a direct comparison of the corresponding binary vectors. To this end, a natural idea is to count the number of coincident spikes in the two vectors. This requires only a logical AND operation per entry to detect two spikes in the same position [67, 66, 1]. This metric is extremely cheap to evaluate, and it intuitively captures a notion of similarity encoded in the timings of the spikes in the two vectors. An alternative metric counts the number of coincident binary digits – thus including not only ones but also zeros. This metric requires an XNOR operation per entry of the encoded time series [61]. Figure 8 illustrates an example for spiking encodings corresponding to a key vector and a query vector.
  • Stochastic computing: The work [55] leverages stochastic computing [2] to evaluate the inner products required by self-attention. Unlike the LIF-based encoding schemes used by the previous class of methods, stochastic computing recovers exact results as the virtual time index $t$ grows large. To this end, queries, keys, and values are transformed into binary signals via Bernoulli encoding (recall (12)). This representation enables multiplications to be performed using simple AND gates. In fact, the result of the AND of two independent Bernoulli random variables $\text{Bern}(p_{1})$ and $\text{Bern}(p_{2})$ is a Bernoulli random variable $\text{Bern}(p_{1}p_{2})$. Therefore, by the law of large numbers, as the virtual time index increases, the estimated rate of spike coincidences detected by the AND gates tends to the correct product value $p_{1}p_{2}$.